### v1

# Shannon Capacity of Channels with Markov Insertions, Deletions and Substitutions

Abstract
We consider channels with synchronization errors modeled as insertions and deletions. A classical result for such channels is the information stability of such channels, hence the existence of the Shannon capacity, when the synchronization errors are memoryless. In this paper, we extend this result to the case where the insertions and deletions have memory. Specifically, we assume that the synchronization errors are governed by a stationary and ergodic finite state Markov chain, and prove that mutual information capacity of such channels exist, and it is equal to its coding capacity, showing that there exists a coding scheme which achieves this limit.
1 Introduction
A mathematical model of a physical channel is defined by conditional distributions $W(y|x)$ for all $x\in{\mathcal{X}},y\in{\mathcal{Y}}$, which means that if a symbol $x\in{\mathcal{X}}$ is input, output $y\in{\mathcal{Y}}$ is observed with probability [^1] This is valid only for a countable ${\mathcal{Y}}$; for uncountable case, there should be output PDFs instead; in this paper uncountable ${\mathcal{Y}}$ are not considered. $W(y|x)$. The fact that $W$ is a channel with the input set ${\mathcal{X}}$ and the output set ${\mathcal{Y}}$ will be denoted by $W:{\mathcal{X}}\rightsquigarrow{\mathcal{Y}}$. The famous Shannon’s theorem [1] considers a memoryless channel $W:{\mathcal{X}}\rightsquigarrow{\mathcal{Y}}$ that is used $n$ times independently. The equivalent channel that transmits $n$ symbols simultaneously is the channel $W^{n}:{\mathcal{X}}^{n}\rightsquigarrow{\mathcal{Y}}^{n}$, where the conditional probabilities are given by $W^{n}(y_{1}^{n}|x_{1}^{n})=\prod_{i=1}^{n}W(y_{i}|x_{i})$, where we use vector notation $a_{b}^{c}=(a_{b},a_{b+1},\dots,a_{c})$.
Consider now an insertion channel $V_{n}$ with $n$ input symbols $x\in{\mathcal{X}}^{n}$. In general, for each input symbol $x_{i}\in{\mathcal{X}}$, an output vector $y_{i}$ over ${\mathcal{X}}$ of length $\geq 1$ is produced. In contrast with the channel $W^{n}$, the resulting output of $V_{n}$ is a concatenation of output vectors, but not the vector $y_{1}^{n}$ of output vectors. For example, if input bits are $(1,0,0)$ and the corresponding outputs for each bit are $(1,0)$, $(0)$ and $(1,0)$, then the overall channel output would be $(1,0,0,1,0)$, not $((1,0),(0),(1,0))$.
In general, a communication setup, besides conditional probabilities of the original channel, also includes some underlying method of combining the outputs of multiple channel uses. In the case of simple memoryless channels, the combining method is just stacking together output symbols into the output vector; in the case of insertion channel, the combining method is concatenation.
In general we can say that the transition probabilities for multiple uses of a channel are not uniquely defined by the conditional probabilities of the single-use (or “original”) channel. For this reason, in the non-trivial setup, instead of a channel, a channel sequence $(W_{n}:{\mathcal{X}}_{n}\rightsquigarrow{\mathcal{Y}}_{n})_{n\in{\mathbb{N}}}$ is usually considered, where sets ${\mathcal{X}}_{n}$, ${\mathcal{Y}}_{n}$ can vary for different $n$. For simplicity, we consider the case when input set for $W_{n}$ is ${\mathcal{X}}_{n}={\mathcal{X}}^{n}$ and ${\mathcal{X}}$ is fixed for all $n$.
For a channel sequence the information capacity can be defined as $\lim_{n\to\infty}{\mathbf{I}}(W_{n})/n$, where ${\mathbf{I}}(W_{n})$ denotes the maximum mutual information between the input and the output of a channel $W_{n}$ over all possible input distributions. The coding capacity (or operational capacity) is the (maximum) asymptotic rate of coding schemes which achieve arbitrary small error probability as $n\to\infty$. The capacity theorem is a statement that the information capacity is equal to coding capacity.
There are many capacity theorems that generalize the Shannon’s theorem on non-trivial channel sequences. Dobrushin’s work [2] considers the most general case, where input and output alphabets can be continuous, and the metric of closeness of decoder’s output to the original signal can be arbitrary . It is shown that the so-called “information stability” of a channel sequence is sufficient for the capacity theorem to hold. In [3], Dobrushin proves the capacity theorem for channels with independent synchronization errors. In that model, for each input symbol, an output vector of finite expected length is produced, and all the output vectors are concatenated to form the resulting output. Since this channel model is general and includes any insertion/deletion/substitution errors, we call such model an IDS channel. The generalization of the capacity theorem to the case of IDS channels with continuous input alphabet is studied in [4]. Capacity theorems are also proven for finite-state Markov channels (FSMC) [5], for Gilbert-Elliott channels [6], for deletion channel concatenated with FSMC [7].
IDS channels are used to model different communication channels, including bit patterned media recording [8] and DNA storage channels [9]. For the case of memoryless insertion/deletion channels, capacity upper bounds and lower bounds are also obtained in [10, 11, 12], and the asymptotic expansion of the i.i.d. deletion channel capacity is found in [13]. See [14, 15] for a detailed review of progress in this area. While most existing results focus on the scenario with independent synchronization errors, it is also important to study the case with insertions/deletions with memory. For instance, current reading and writing technologies in DNA sequencing introduce memory to the channel errors [16, 17]. A simple and useful model for synchronization errors with memory is a Markov chain.
In this paper, we consider IDS channels where the underlying insertions/deletions are governed by a Markov chain, referred to as a Markov-IDS channel. That is, each state of the Markov chain is an IDS channel, called component channel, and the overall output of Markov-IDS channel is obtained by concatenating the individual component channel outputs. The Markov-IDS channel is a generalization of both IDS and FSMC: it is more general than IDS since it changes over time according to Markov chain rule; it is more general then FSMC since it concatenates the outputs of each component IDS channel (which can be translated to non-concatenated version via inserting a special symbol at the end of each component channel’s output). We note that while this model generalizes the synchronization errors as studied by Dobrushin, it is not sufficiently general to encompass some DNA channel models with insertions/deletions with memory as in [16, 17] since the channel states do not depend on the channel input in our model. Further extensions are left for future work.
We also introduce some unified notation, which helps to elucidate the idea behind Dobrushin’s methods, which can be seen as applying functions to channel sequences. Namely, if a function is applied on a channel’s output, and this function is not “doing much” in a certain sense, then it changes neither the mutual information capacity nor the coding capacity of the original channel. This means that one can apply any finite chain of such functions and still have the same capacities.
The paper is organized as follows. In Section 2, notations and information-theoretic definitions are introduced. In Section 3, functions on channels and their properties are described. In Section 4, the Markov-IDS channel sequence is defined. In Section 5, the capacity theorem for Markov-IDS channel sequence is proven. In Section 6, the conclusion is drawn.
2 Background
2.1 Notations
We use the following notations. Sets are written in calligraphic letters ${\mathcal{X}},{\mathcal{Y}},{\overline{\mathcal{Y}}}$; operators are written in bold letters ${\mathbf{C}},{\mathbf{E}},{\mathbf{I}}$. For $b,c\in{\mathbb{N}}$, we denote a vector $a_{b}^{c}=(a_{b},a_{b+1},\dots,a_{c})$. Here, elements $a_{i}$ can be of any type: a number, a distribution, a channel, etc. Also, we denote infinite sequences as $a^{\infty}=(a_{1},a_{2},\dots)$. For a set ${\mathcal{A}}$, we denote a set of all finite sequences with elements from ${\mathcal{A}}$ by ${\overline{\mathcal{A}}}=\left\{{()}\right\}\cup\bigcup_{i=1}^{\infty}{\mathcal{A}}^{i}.$ For a vector $a=a_{1}^{n}\in{\overline{\mathcal{A}}}$ we denote by ${\mathbf{N}}(a)=n$ the vector length. We denote the concatenation (gluing) operator $\mathbf{g}$ as $\mathbf{g}(a_{1}^{n},b_{1}^{m})=(a_{1},a_{2},...,a_{n},b_{1},...,b_{m}),$ which is also generalized onto any number of input vectors. Also, define the function $\mathbf{g}_{n}:{\overline{\mathcal{A}}}^{n}\to{\overline{\mathcal{A}}}$, which glues $n$ vectors ${\overline{a}}_{i}\in{\overline{\mathcal{A}}}$ as $\mathbf{g}_{n}({\overline{a}}_{1}^{n})=\mathbf{g}({\overline{a}}_{1},{\overline{a}}_{2},...,{\overline{a}}_{n}).$ The inverse image $\mathbf{g}^{-1}_{n}(a)$ is the set of all possible ways of cutting $a\in{\overline{\mathcal{A}}}$ into $n$ subvectors, i.e., $\mathbf{g}^{-1}_{n}(a)=\left\{{{\overline{a}}_{1}^{n}|\mathbf{g}_{n}({\overline{a}}_{1}^{n})=a}\right\}$. For a set ${\mathcal{S}}$, we denote by ${\mathbb{D}}_{\mathcal{S}}$ the set of all distributions over ${\mathcal{S}}$, i.e.
${\mathbb{D}}_{\mathcal{S}}=\left\{{f:{\mathcal{S}}\to[0,1]\left|\sum_{x\in{\mathcal{S}}}f(x)=1\right.}\right\}.$
For $n\in{\mathbb{N}}$, denote by ${\mathbb{D}}_{n}$ the set of distributions over the set $\left\{{1,2,\dots,n}\right\}$ in a form of row vectors $\rho_{1}^{n}$, $\sum_{i=1}^{n}\rho_{i}=1$, $\rho_{i}\geq 0$.
In general, probabilities are usually defined using sigma-algebras. In this paper, however, all probabilities are defined on the countable sets, so the probability measure is always completely defined by the probabilities of the elementary events $\Pr{\left[{X=x}\right]}$.
The probability distribution for any probability, expectation or entropy is written as a subscript. We assume summation of all values of random variables, which are not fixed. For example, if we have a conditional distribution $W(u,x|v,y)$, where $u\in{\mathcal{U}},v\in{\mathcal{V}},x\in{\mathcal{X}},y\in{\mathcal{Y}}$, and some specific values of $u$ and $v$ (or fixed them by $\exists$,$\forall$ , ∀ symbols). Then,
$\Pr_{W(u,x|v,y)}{\left[{x>0}\right]}$ $=\sum_{\begin{subarray}{c}x\in{\mathcal{X}}:x>0\\ y\in{\mathcal{Y}}\end{subarray}}W(u,x|v,y),$
${\mathbf{E}}_{W(u,x|v,y)}{\left[{f(x,y)}\right]}$ $=\sum_{x\in{\mathcal{X}},y\in{\mathcal{Y}}}W(u,x|v,y)\cdot f(x,y).$
Entropy of $x$ is expectation of $-\log_{2}x$, so ${\mathbf{H}}_{A}{\left[{B}\right]}={\mathbf{E}}_{A}{\left[{-\log_{2}B}\right]}$. If $A=B$, then we omit the subscript for simplicity:
${\mathbf{H}}_{W(x,y)}{\left[{W(x|y)}\right]}$ $=-\sum_{x\in{\mathcal{X}},y\in{\mathcal{Y}}}W(x,y)\cdot\log_{2}W(x|y),$
${\mathbf{H}}{\left[{W(x,y)}\right]}$ $=-\sum_{x\in{\mathcal{X}},y\in{\mathcal{Y}}}W(x,y)\cdot\log_{2}W(x,y).$
2.2 Channels and channel sequences
We define a channel $W$ as a collection of conditional distributions $W(y|x)$, defined $\forall x\in{\mathcal{X}}$, $\forall y\in{\mathcal{Y}}$. Throughout the paper, we assume a finite set ${\mathcal{X}}$ and a countable set ${\mathcal{Y}}$. Formally, a channel $W$ is a function $W:{\mathcal{X}}\times{\mathcal{Y}}\to[0,1]$, such that
$\forall x\in{\mathcal{X}}:\sum_{y\in{\mathcal{Y}}}W(y|x)=1.$
The fact that $W$ has input messages from ${\mathcal{X}}$ and output from ${\mathcal{Y}}$ is denoted by $W:{\mathcal{X}}\rightsquigarrow{\mathcal{Y}}$.
A channel sequence $W$ is defined as an infinite sequence of channels
$W=(W_{n})_{n\in{\mathbb{N}}},\;W_{i}:{\mathcal{X}}^{i}\rightsquigarrow{\mathcal{Y}}_{i}\;\forall i\in{\mathbb{N}}.$
Note that for any $i$, the output set ${\mathcal{Y}}_{i}$ is arbitrary, but the input set for the $i$-th channel is ${\mathcal{X}}^{i}$, so the $i$-th channel transmits exactly $i$ symbols from ${\mathcal{X}}$.
Channel $W$ defines the conditional distribution of the output given its input. If we further define some input distribution $p\in{\mathbb{D}}_{\mathcal{X}}$, then we can compute the joint distribution of input and output, which is denoted by $p\circ W$:
$p\circ W(x,y)=p(x)W(y|x)$ (1a)
$p\circ W(*,y)=\sum_{x\in{\mathcal{X}}}p(x)W(y|x),$ (1b)
where (1b) defines the probability of output $y$, given the input distribution $p$.
For channels $W_{i}:{\mathcal{X}}_{i}\rightsquigarrow{\mathcal{Y}}_{i}$, define the channel product $W=\prod_{i=1}^{n}W_{i}$ as a channel $W:{\mathcal{X}}_{1}\times\dots\times{\mathcal{X}}_{n}\rightsquigarrow{\mathcal{Y}}_{1}\times\dots\times{\mathcal{Y}}_{n}$, such that
$W=\prod_{i=1}^{n}W_{i}\iff W(y_{1}^{n}|x_{1}^{n})=\prod_{i=1}^{n}W_{i}(y_{i}|x_{i})$ (2)
A similar notation is used for the $n$-th power $W^{n}:{\mathcal{X}}^{n}\rightsquigarrow{\mathcal{Y}}^{n}$ of a channel $W:{\mathcal{X}}\rightsquigarrow{\mathcal{Y}}$.
2.3 Mutual information capacity
The mutual information density of channel $W$ for input distribution $p\in{\mathbb{D}}_{\mathcal{X}}$ and given values of input $x\in{\mathcal{X}}$ and output $y\in{\mathcal{Y}}$ is defined as:
$\mathbf{i}(W,p,x,y)\stackrel{{\scriptstyle\triangle}}{{=}}\begin{cases}0,\text{if }W(y|x)=0\text{ or }p\circ W(*,y)=0\\ \log_{2}\frac{W(y|x)}{p\circ W(*,y)},\text{ otherwise}\end{cases}$ (3)
where $p\circ W(*,y)$ is given by (1b).The mutual information of $W$ under input distribution $p$ is the expectation of mutual information density:
${\mathbf{I}}(W,p)$ $\stackrel{{\scriptstyle\triangle}}{{=}}{\mathbf{E}}_{p\circ W(x,y)}[\mathbf{i}(W,p,x,y)]=\sum_{x\in{\mathcal{X}},y\in{\mathcal{Y}}}p\!\circ\!W(x,y)\cdot\mathbf{i}(W,p,x,y).$ (4)
The mutual information capacity/i-capacity [2] of channel $W$ is the maximum possible mutual information of $W$ under any input distribution, and is given by [^2] In [2], the setup is for the much more general case, and “$\sup$” over the input distributions is used instead of “$\max$”. In our setup we assumed (see Section 2.2) that ${\mathcal{X}}$ is finite and ${\mathcal{Y}}$ is countable, so the maximum is always achieved as a maximum of a continuous function over a compact set.
${\mathbf{I}}(W)\stackrel{{\scriptstyle\triangle}}{{=}}\max_{p\in{\mathbb{D}}_{\mathcal{X}}}{\mathbf{I}}(W,p).$ (5)
We will call the distribution $p$, for which the maximum is achieved, as the optimal distribution (for $W$). By the mutual information capacity/i-capacity of channel sequence $V=V^{\infty}$ we mean the asymptotic average i-capacity of $V_{n}$ per input symbol:
${\mathbf{I}}(V)\stackrel{{\scriptstyle\triangle}}{{=}}\lim_{n\to\infty}\frac{{\mathbf{I}}(V_{n})}{n},$ (6)
if the limit exists. Note that the i-capacity of a channel always exists, since ${\mathbf{I}}(W,p)$ is a concave function of the input distribution $p$ with countable number of coefficients; however, the i-capacity of a channel sequence might not exist, since sequence ${\mathbf{I}}(V_{n})/n$ might not converge [^3] Consider a channel which is ideal for even $n$ and complete noise for odd $n$. In this case, the sequence of values of ${\mathbf{I}}(V_{n})/n$ is $(\log_{2}|{\mathcal{X}}|,0,\log_{2}|{\mathcal{X}}|,0,\dots)$. Such sequence has no limit for $|{\mathcal{X}}|>1$..
2.4 Coding capacity
For a channel $W:{\mathcal{X}}\rightsquigarrow{\mathcal{Y}}$, code size $M\in{\mathbb{N}}$, $M\leq|{\mathcal{X}}|$ and error probability $\varepsilon\in[0,1]$, we define a $(W,M,\varepsilon)$-coding scheme as a collection $(x_{1}^{M},{\mathcal{R}}_{1}^{M}=({\mathcal{R}}_{1},{\mathcal{R}}_{2},\dots,{\mathcal{R}}_{M}))$, where:
  • all $x_{i}$ are distinct codewords from ${\mathcal{X}}$, $i=1\dots M$;
  • ${\mathcal{R}}_{1}^{M}$ is a collection of $M$ disjoint sets ${\mathcal{R}}_{i}\subseteq{\mathcal{Y}}$, which are the decoder’s decision regions. When $y\in{\mathcal{R}}_{i}$ is received, the decoder outputs $x_{i}$;
  • the decoding error probability is not greater than $\varepsilon$ for any $x_{i}$, i.e.
    $\forall i\in\left\{{1,\dots,M}\right\}:\sum_{y\notin{\mathcal{R}}_{i}}{W(y|x_{i})}\leq\varepsilon.$
This definition obviously implies that if there exists a $(W,M,\varepsilon)$ coding scheme, there also exists a $(W,M^{\prime},\varepsilon^{\prime})$-coding scheme for any $M^{\prime}\leq M$ and $\varepsilon^{\prime}\geq\varepsilon$.
Consider a channel sequence $W$, where $W_{n}:{\mathcal{X}}^{n}\rightsquigarrow{\mathcal{Y}}_{n}$. We call $R\in{\mathbb{R}}_{\geq 0}$ an achievable rate over $W$, if
[…]
We denote by ${\mathbf{R}}(W)$ the set of achievable rates $R$ for channel sequence $W$. By the coding capacity/c-capacity ${\mathbf{C}}(W)$ of the channel sequence $W$ we mean the supremum of all achievable rates [3]:
${\mathbf{C}}(W)$ $\stackrel{{\scriptstyle\triangle}}{{=}}\sup{\mathbf{R}}(W)$ (8)
Note that, $0\in{\mathbf{R}}(W)$ for any $W$, and if $R\in{\mathbf{R}}(W)$, then $R^{\prime}\in{\mathbf{R}}(W)$ for any rate $R^{\prime}<R$. Thus, the set ${\mathbf{R}}(W)$ is a connected set, i.e., there always exists $\alpha\in[0,\log_{2}|{\mathcal{X}}|]$, such that ${\mathbf{R}}(W)\in\left\{{[0,\alpha],[0,\alpha)}\right\}$. This implies that ${\mathbf{C}}(W)=\alpha$, i.e., any channel sequence has the c-capacity.
In [18], the coding capacity (c-capacity) ${\mathbf{C}}(W)$ is derived in terms of mutual information density:
${\mathbf{C}}(W)=\sup\left\{{\alpha\left|\lim_{n\to\infty}~{}\min_{p_{n}\in{\mathbb{D}}_{{\mathcal{X}}^{n}}}\Pr_{p_{n}\circ W_{n}(x,y)}\left\{{\frac{\mathbf{i}(W_{n},p_{n},x,y)}{n}\leq\alpha}\right\}=0\right.}\right\}$ (9)
2.5 Information-stable channels
The channel sequence $W$ is called information-stable, if it has i-capacity, i.e., ${\mathbf{I}}(W)$ exists, and for optimal distributions $p^{*}_{n}\in{\mathbb{D}}_{{\mathcal{X}}^{n}}$ for $W_{n}$, for any $\delta>0$:
$\lim_{n\to\infty}~{}\Pr_{p^{*}_{n}\circ W_{n}(x,y)}{\left[{\left|{\frac{\mathbf{i}(W_{n},p^{*}_{n},x,y)}{n}-{\mathbf{I}}(W)}\right|>\delta}\right]}=0.$ (10)
Proposition 1 (Theorem 1 in [2]).
If a channel sequence $W$ is information-stable, then ${\mathbf{C}}(W)={\mathbf{I}}(W)$.
3 Functions of channels and their properties
For a channel $W:{\mathcal{X}}\rightsquigarrow{\mathcal{Y}}$ and function $f:{\mathcal{Y}}\to{\mathcal{Z}}$ we write $W^{\prime}={f}{\left({W}\right)}$ to refer to a channel $W^{\prime}:{\mathcal{X}}\rightsquigarrow{\mathcal{Z}}$, defined as a composition of channel $W$ and function $f$. That is,
$W^{\prime}(z|x)=\sum_{y\in f^{-1}(z)}W(y|x),$ (11)
where $f^{-1}(z)=\left\{{y:f(y)=z}\right\}$ is the inverse image of $f$. Channel $W^{\prime}$ is degraded with respect to $W$, meaning that it can be represented as a pipeline of $W$ and a function $f$ at its output. It is well-known and can be straightforwardly checked that ${\mathbf{C}}(W^{\prime})\leq{\mathbf{C}}(W)$, and if both ${\mathbf{I}}(W)$ and ${\mathbf{I}}(W^{\prime})$ exist, then ${\mathbf{I}}(W^{\prime})\leq{\mathbf{I}}(W)$. Let $\Phi=\log_{2}\max_{z\in{\mathcal{Z}}}\left|{f^{-1}(z)}\right|$ be the maximum log-size of a preimage. Then, the following proposition holds.
Proposition 2 (Equation (4.3) in [3]).
For any input distribution $p\in{\mathbb{D}}_{{\mathcal{X}}}$:
${\mathbf{E}}_{p\circ W(x,y)}\bigg{[}\big{|}\mathbf{i}(W,p,x,y)-\mathbf{i}(W^{\prime},p,x,f(y))\big{|}\bigg{]}\leq\Phi.$ (12)
This proposition can be interpreted as follows. The process of forming the (random) output of channel $W$ can be seen as a pipeline of two steps. First, choose the value of $z$ based on probabilities $W^{\prime}(z|x)$. This is equivalent to channel $W^{\prime}$. Second, randomly choose one value among all $\leq 2^{\Phi}$ possible values of $y\in f^{-1}(y^{\prime})$. The maximum entropy of the second step is achieved when all preimages are equiprobable, namely, the entropy of the second step is at most $\Phi$.
For a channel sequence $U_{n}:{\mathcal{X}}^{n}\rightsquigarrow{\mathcal{Y}}_{n}$ and sequence of functions $f_{n}:{\mathcal{Y}}_{n}\to{\mathcal{Z}}_{n}$, we denote by $V={f}{\left({U}\right)}$ a channel sequence $V_{n}={f_{n}}{\left({U_{n}}\right)}:{\mathcal{X}}^{n}\rightsquigarrow{\mathcal{Z}}_{n}$, $n\in{\mathbb{N}}$. Also, we denote by $\Phi_{n}=\log_{2}\max_{z\in{\mathcal{Z}}_{n}}\left|{f_{n}^{-1}(z)}\right|$.
Proposition 3 (Function with small preimage).
If $\Phi_{n}\in o(n)$, then ${\mathbf{C}}(U)={\mathbf{C}}(V)$. If, furthermore, at least one of ${\mathbf{I}}(U)$, ${\mathbf{I}}(V)$ exist, then they both exist and ${\mathbf{I}}(U)={\mathbf{I}}(V)$.
Proof.
First, note that ${\mathbf{E}}_{p_{n}\circ V_{n}}{\left[{\mathbf{i}(V_{n},p_{n},x,y)}\right]}={\mathbf{E}}_{p_{n}\circ U_{n}}{\left[{\mathbf{i}(V_{n},p_{n},x,f(y))}\right]}$. Using Proposition 2, for any $p_{n}$:
${\mathbf{I}}(U_{n},p_{n})-{\mathbf{I}}(V_{n},p_{n})={\mathbf{E}}_{p_{n}\circ U_{n}}{\left[{\mathbf{i}(U_{n},p_{n},x,y)-\mathbf{i}(V_{n},p_{n},x,f_{n}(y))}\right]}$
$\leq{\mathbf{E}}_{p_{n}\circ U_{n}}{\left[{\left|{\mathbf{i}(U_{n},p_{n},x,y)-\mathbf{i}(V_{n},p_{n},x,f_{n}(y))}\right|}\right]}\leq\Phi_{n}.$
In particular, for any sequence of input distributions $\left\{{p_{n}\in{\mathbb{D}}_{{\mathcal{X}}^{n}}}\right\}_{n\in{\mathbb{N}}}$,
$0\displaystyle 0\leq\lim_{n\to\infty}{\mathbf{E}}_{p_{n}\circ U_{n}}{\left[{\left|{\frac{\mathbf{i}(U_{n},p_{n},x,y)-\mathbf{i}(V_{n},p_{n},x,f_{n}(y))}{n}}\right|}\right]}\ \leq\lim_{n\to\infty}\frac{\Phi_{n}}{n}=$ 0 . (13)
To prove that ${\mathbf{C}}(U)={\mathbf{C}}(V)$, we will use the expression of c-capacity given by (9). Assume that ${\mathbf{C}}(U)=\alpha$ and ${\mathbf{C}}(V)\neq\alpha$. Then, ${\mathbf{C}}(V)<\alpha$ and
$\forall\varepsilon>0,\delta>0,\exists n^{*}_{0}:\forall n\geq n_{0},\exists p^{*}:~{}\Pr_{p^{*}\circ U_{n}(x,y)}{\left[{\frac{\mathbf{i}(U_{n},p^{*},x,y)}{n}>\alpha-\varepsilon}\right]}$ $>1-\delta.$ (14)
$\exists\varepsilon^{*}>0,\delta^{*}>0:\forall n_{0},\exists n^{*}\geq n_{0}:\forall p,~{}\Pr_{p\circ U_{n^{*}}(x,y)}{\left[{\frac{\mathbf{i}(V_{n^{*}},p,x,f_{n^{*}}(y))}{n^{*}}\leq\alpha-\varepsilon^{*}}\right]}$ $\geq\delta^{*}.$ (15)
Since (14) holds for any $\varepsilon>0$, $\delta>0$, we can let $\varepsilon=\varepsilon^{*}/2$, $\delta=\delta^{*}/2$, where $\varepsilon^{*}$ and $\delta^{*}$ are from (15), and obtain simultaneously
$\exists\varepsilon^{*}>0,\delta^{*}>0:\begin{cases}\exists n^{*}_{0}:\forall n\geq n^{*}_{0},\exists p_{n}^{*}:~{}\Pr_{p_{n}^{*}\circ U_{n}(x,y)}{\left[{\mathbf{i}(U_{n},p_{n}^{*},x,y)/n>\alpha-\varepsilon^{*}/2}\right]}>1-\delta^{*}/2\\ \forall n_{0},\exists n^{*}\geq n_{0}:\forall p_{n^{*}},~{}\Pr_{p_{n^{*}}\circ U_{n^{*}}(x,y)}{\left[{\mathbf{i}(V_{n^{*}},p_{n^{*}},x,f_{n^{*}}(y))/{n^{*}}\leq\alpha-\varepsilon^{*}}\right]}\geq\delta^{*}\end{cases}.$
So, the first inequality holds for all $n$ starting from $n^{*}_{0}$, and for each $n$ specific $p_{n}^{*}$ should be picked up. The second inequality holds for some $n$ larger than arbitrary $n_{0}$, for all $p_{n}$. Bringing these two conditions together, we can satisfy both inequalities by doing as follows:
  • for any $n_{0}$, pick up the specific value of $n^{*}\geq\max\left\{{n_{0},n^{*}_{0}}\right\}$, for which the second inequality holds;
  • since $n^{*}\geq n^{*}_{0}$, the first equality still holds for some specific value of $p^{*}_{n^{*}}$. The second inequality will hold for $p_{n^{*}}^{*}$ as well, since it holds for any $p_{n^{*}}$.
The above procedure results in
$*.\displaystyle\exists\varepsilon^{*}>0,\delta^{*}>0:\forall n_{0},\exists n\geq n_{0},\exists p_{n}:\begin{cases}\Pr_{p_{n}\circ U_{n}(x,y)}{\left[{\mathbf{i}(U_{n},p_{n},x,y)/n>\alpha-\varepsilon^{*}/2}\right]}>1-\delta^{*}/2\\ \Pr_{p_{n}\circ U_{n}(x,y)}{\left[{\mathbf{i}(V_{n},p_{n},x,f_{n}(y))/{n}\leq\alpha-\varepsilon^{*}}\right]}\geq\delta^{*}\end{cases}.$
Note that both probabilities are defined over the same probability space. The sum of the two probabilities is greater than $21+\delta^{*}/$ 2, which means that with probability greater than $\delta^{*}/2$ the difference $\mathbf{i}(U_{n},p_{n},x,y)/n-\mathbf{i}(V_{n},p_{n},x,f_{n}(y))/{n}$ is greater than $\varepsilon^{*}/2$, which implies
$\exists\varepsilon^{*}>0,\delta^{*}>0:\forall n_{0},\exists n\geq n_{0},\exists p_{n}:{\mathbf{E}}_{p_{n}\circ U_{n}}{\left[{\left|{\frac{\mathbf{i}(U_{n},p_{n},x,y)-\mathbf{i}(V_{n},p_{n},x,f_{n}(y))}{n}}\right|}\right]}>\frac{\varepsilon^{*}\delta^{*}}{4},$
which contradicts (13).
Equality of i-capacities follows from (13). Let $p^{*}_{n}$ be the optimal distribution for $U_{n}$. If either ${\mathbf{I}}(U)$ or ${\mathbf{I}}(V)$ exists, then
$0\leq{\mathbf{I}}(U)-{\mathbf{I}}(V)\leq\lim_{n\to\infty}\frac{{\mathbf{I}}(U_{n},p^{*}_{n})-{\mathbf{I}}(V_{n},p^{*}_{n})}{n}\leq\lim_{n\to\infty}\frac{\Phi_{n}}{n}=0,$
so ${\mathbf{I}}(U)={\mathbf{I}}(V)$. ∎
Thus, one can apply such function sequences to a channel sequence $W$, changing neither ${\mathbf{I}}(W)$ nor ${\mathbf{C}}(W)$.
Proposition 4 (Semi-bijective function).
Consider channels $U:{\mathcal{X}}\rightsquigarrow{\mathcal{Y}}$ and $V:{\mathcal{X}}\rightsquigarrow{\mathcal{Z}}$, such that $V=f(U)$, $f:{\mathcal{Y}}\to{\mathcal{Z}}$. Consider partitions ${\mathcal{Y}}={\mathcal{A}}\sqcup{\mathcal{B}}$, ${\mathcal{Z}}={\mathcal{A}}^{\prime}\sqcup{\mathcal{B}}^{\prime}$, such that $f({\mathcal{A}})={\mathcal{A}}^{\prime}$ and $f$ is bijective between ${\mathcal{A}}$ and ${\mathcal{A}}^{\prime}$. Denote by $\beta(x)=\Pr_{U(y|x)}{\left[{y\in{\mathcal{B}}}\right]}$ the probability that the output of $U$ belongs to ${\mathcal{B}}$. If $\forall x:\beta(x)\leq\overline{\beta}$, then
${\mathbf{I}}(U)-{\mathbf{I}}(V)\leq\overline{\beta}\log_{2}|{\mathcal{X}}|+\frac{1}{e\ln 2}.$ (16)
Proof.
Let $U^{\prime}:{\mathcal{X}}\rightsquigarrow{\mathcal{B}}$ be a channel equivalent to $U$ with an additional condition that the output belongs to ${\mathcal{B}}$, with the normalized transition probabilities
$U^{\prime}(y|x)=\frac{U(y|x)}{\beta(x)}.$
Consider some $p\in{\mathbb{D}}_{\mathcal{X}}$. By definition, ${\mathbf{I}}(U,p)={\mathbf{E}}_{p\circ U(x,y)}{\left[{\mathbf{i}(U,p,x,y)}\right]}$. The summation, corresponding to taking the expectation, can be expressed by independent summation over ${\mathcal{A}}$ and ${\mathcal{B}}$:
${\mathbf{E}}_{p\circ U(x,y)}{\left[{\mathbf{i}(U,p,x,y)}\right]}=\sum_{y\in{\mathcal{Y}}}\sum_{x\in{\mathcal{X}}^{n}}p(x)U(y|x)\mathbf{i}(U,p,x,y)$
$=\underbrace{\sum_{y\in{\mathcal{A}}}\sum_{x\in{\mathcal{X}}}p(x)U(y|x)\mathbf{i}(U,p,x,y)}_{a}+\underbrace{\sum_{y\in{\mathcal{B}}}\sum_{x\in{\mathcal{X}}}p(x)U(y|x)\mathbf{i}(U,p,x,y)}_{b}$
Let us analyze term $b$. First, introduce modified distribution ${\widetilde{p}}(x)=p(x)\beta(x)/\Delta$, where $\Delta=\sum_{x}p(x)\beta(x)$. Note that $0\leq\Delta\leq\overline{\beta}$. Substituting $p(x)U(y|x)=\Delta{\widetilde{p}}(x)U^{\prime}(y|x)$ for $y\in{\mathcal{B}}$, one obtains
$b=\sum_{y\in{\mathcal{B}},x\in{\mathcal{X}}}p(x)U(y|x)\log_{2}\frac{U(y|x)}{\sum_{w}p(w)U(y|w)}=\Delta\sum_{y\in{\mathcal{B}},x\in{\mathcal{X}}}{\widetilde{p}}(x)U^{\prime}(y|x)\log_{2}\frac{\beta(x)U^{\prime}(y|x)}{\Delta\sum_{w}{\widetilde{p}}(w)U^{\prime}(y|w)}$
$=\Delta\cdot{\mathbf{I}}(U^{\prime},{\widetilde{p}})+\Delta\sum_{y\in{\mathcal{B}},x\in{\mathcal{X}}}{\widetilde{p}}(x)U^{\prime}(y|x)\log_{2}\frac{\beta(x)}{\Delta}$
$=\Delta{\left({{\mathbf{I}}(U^{\prime},{\widetilde{p}})-\log_{2}\Delta}\right)}+\sum_{y\in{\mathcal{B}},x\in{\mathcal{X}}}p(x)U^{\prime}(y|x)\beta(x)\log_{2}\beta(x)$
Using inequality $-\frac{1}{e\ln 2}\leq x\log_{2}x\leq 0$ when $0<x\leq 1$, one obtains
$-\Delta\log_{2}\Delta-\frac{1}{e\ln 2}\leq b\leq\Delta\log_{2}\frac{|{\mathcal{X}}|}{\Delta}$ (17)
The same can be done for the channel $V$:
${\mathbf{E}}_{p\circ V(x,y)}{\left[{\mathbf{i}(V,p,x,y)}\right]}=\sum_{y\in{\mathcal{A}}^{\prime}}\sum_{x\in{\mathcal{X}}}p(x)V(y|x)\mathbf{i}(V,p,x,y)+\underbrace{\sum_{y\in{\mathcal{B}}^{\prime}}\sum_{x\in{\mathcal{X}}}p(x)V(y|x)\mathbf{i}(V,p,x,y)}_{b^{\prime}}$
$=\sum_{y\in{\mathcal{A}}}\sum_{x\in{\mathcal{X}}}p(x)U(y|x)\mathbf{i}(U,p,x,y)+b^{\prime}=a+b^{\prime}$
Note that $\beta(x)=\sum_{y\in{\mathcal{B}}}U(y|x)=\sum_{z\in{\mathcal{B}}^{\prime}}V(z|x)$, so, using the auxiliary channel $V^{\prime}(z|x):{\mathcal{X}}\rightsquigarrow{\mathcal{B}}^{\prime}$ defined as $V^{\prime}(z|x)=V(z|x)/\beta(x)$, one obtains for any $z\in{\mathcal{B}}^{\prime}$, $p(x)V(z|x)=\Delta{\widetilde{p}}(x)V^{\prime}(z|x)$. Similarly to $b$, the value of $b^{\prime}$ can be represented as
$b^{\prime}=\sum_{z\in{\mathcal{B}}^{\prime},x\in{\mathcal{X}}}p(x)V(z|x)\mathbf{i}(V,p,x,z)=\Delta\sum_{z\in{\mathcal{B}}^{\prime},x\in{\mathcal{X}}}{\widetilde{p}}(x)V^{\prime}(z|x)\log_{2}\frac{\beta(x)V^{\prime}(z|x)}{\Delta\sum_{w}{\widetilde{p}}(w)V^{\prime}(z|w)}$
$=\Delta{\left({{\mathbf{I}}(V^{\prime},{\widetilde{p}})-\log_{2}\Delta}\right)}+\sum_{z\in{\mathcal{B}}^{\prime},x\in{\mathcal{X}}}p(x)V^{\prime}(y|x)\beta(x)\log_{2}\beta(x)$
so (17) holds for $b^{\prime}$ as well. This implies
$\forall p\in{\mathbb{D}}_{\mathcal{X}}:{\mathbf{I}}(U,p)-{\mathbf{I}}(V,p)=b-b^{\prime}\leq\Delta\log_{2}\frac{|{\mathcal{X}}|}{\Delta}+\Delta\log_{2}\Delta+\frac{1}{e\ln 2}=\Delta\log_{2}|{\mathcal{X}}|+\frac{1}{e\ln 2}$
Since $\Delta\leq\overline{\beta}$, (16) holds. ∎
Proposition 5 (Function that does nothing almost surely).
Consider channel sequences $U_{n}:{\mathcal{X}}^{n}\rightsquigarrow{\mathcal{Y}}_{n}$ and $V_{n}:{\mathcal{X}}^{n}\rightsquigarrow{\mathcal{Z}}_{n}$, such that $V_{n}=f_{n}(U_{n})$, $f_{n}:{\mathcal{Y}}_{n}\to{\mathcal{Z}}_{n}$. Consider partitions ${\mathcal{Y}}_{n}={\mathcal{A}}_{n}\sqcup{\mathcal{B}}_{n}$, ${\mathcal{Z}}_{n}={\mathcal{A}}^{\prime}_{n}\sqcup{\mathcal{B}}^{\prime}_{n}$, such that $f_{n}({\mathcal{A}}_{n})={\mathcal{A}}^{\prime}_{n}$ and $f_{n}$ is bijective between ${\mathcal{A}}_{n}$ and ${\mathcal{A}}^{\prime}_{n}$. Denote by $\beta_{n}(x)=\Pr_{U_{n}(y|x)}{\left[{y\in{\mathcal{B}}_{n}}\right]}$, and $\overline{\beta}_{n}=\max_{x}\beta_{n}(x)$. If the output of channel $U_{n}$ almost always belongs to ${\mathcal{A}}_{n}$, i.e., $\lim_{n\to\infty}\overline{\beta}_{n}=0,$ then ${\mathbf{C}}(U)={\mathbf{C}}(V)$. If, in addition, one of ${\mathbf{I}}(U)$, ${\mathbf{I}}(V)$ exists, then they both exist and ${\mathbf{I}}(U)={\mathbf{I}}(V)$.
Proof.
If either ${\mathbf{I}}(U)$ or ${\mathbf{I}}(V)$ exists, then, using proposition 4,
${\mathbf{I}}(U)-{\mathbf{I}}(V)=\lim_{n\to\infty}\frac{{\mathbf{I}}(U_{n})-{\mathbf{I}}(V_{n})}{n}=\lim_{n\to\infty}{\left({\frac{\overline{\beta}_{n}n\log_{2}|{\mathcal{X}}|}{n}+\frac{1}{ne\ln 2}}\right)}=0,$
and both ${\mathbf{I}}(U)$ and ${\mathbf{I}}(V)$ exist.
According to (9), proving that ${\mathbf{C}}(U)={\mathbf{C}}(V)$ can be done analyzing the probability of $\mathbf{i}(V_{n},p_{n},x,y)$ being less than $\alpha n$ for some value of $\alpha$. If $\mathbf{i}(V_{n},p_{n},x,y)=\mathbf{i}(U_{n},p_{n},x,y)$ and $\mathbf{i}(U_{n},p_{n},x,y)>\alpha n$, then surely $\mathbf{i}(V_{n},p_{n},x,y)>\alpha n$. This can be used as a lower bound on probability:
$\Pr_{p_{n}\circ V_{n}(x,y)}{\left[{\mathbf{i}(V_{n},p_{n},x,y)\leq\alpha n}\right]}=\Pr_{p_{n}\circ U_{n}(x,y)}{\left[{\mathbf{i}(V_{n},p_{n},x,f(y))\leq\alpha n}\right]}$
$\leq\Pr_{p_{n}\circ U_{n}(x,y)}{\left[{\mathbf{i}(V_{n},p_{n},x,f(y))\neq\mathbf{i}(U_{n},p_{n},x,y)\vee\mathbf{i}(V_{n},p_{n},x,f(y))=\mathbf{i}(U_{n},p_{n},x,y)\leq\alpha n}\right]}$
$\leq\Pr_{p_{n}\circ U_{n}(x,y)}{\left[{\mathbf{i}(V_{n},p_{n},x,f(y))\neq\mathbf{i}(U_{n},p_{n},x,y)\vee\mathbf{i}(U_{n},p_{n},x,y)\leq\alpha n}\right]}$
$\leq\Pr_{p_{n}\circ U_{n}(x,y)}{\left[{\mathbf{i}(V_{n},p_{n},x,f(y))\neq\mathbf{i}(U_{n},p_{n},x,y)}\right]}+\Pr_{p_{n}\circ U_{n}(x,y)}{\left[{\mathbf{i}(U_{n},p_{n},x,y)\leq\alpha n}\right]}$
In the last expression, with $n\to\infty$, the first probability goes to zero for any $p_{n}$. So,
$\lim_{n\to\infty}\min_{p_{n}}\Pr_{p_{n}\circ U_{n}(x,y)}{\left[{\mathbf{i}(U_{n},p_{n},x,y)\leq\alpha n}\right]}=0\implies\lim_{n\to\infty}\min_{p_{n}}\Pr_{p_{n}\circ V_{n}(x,y)}{\left[{\mathbf{i}(V_{n},p_{n},x,y)\leq\alpha n}\right]}=0,$
which means that the set of such $\alpha$’s for $U_{n}$ is a subset of such set for $V_{n}$, which implies ${\mathbf{C}}(V)\geq{\mathbf{C}}(U)$ and thus ${\mathbf{C}}(V)={\mathbf{C}}(U)$.
If this proposition is combined with Proposition 3, the result is that one can apply any finite number of functions, which either have sufficiently small pre-image or sufficiently small probability of changing the output, without changing neither ${\mathbf{I}}$ nor ${\mathbf{C}}$ of the original channel sequence.
4 The Markov-IDS channel sequence
4.1 The IDS channel sequence
In [3] a general channel model, which represents any channel with independent insertions, deletions, and substitutions is considered.
Consider a basic channel $W:{\mathcal{X}}\rightsquigarrow{\overline{\mathcal{Y}}}$, such that for each $x\in{\mathcal{X}}$ the expected length ${\mathbf{N}}(y)$ of the output sequence $y\in{\overline{\mathcal{Y}}}$ is finite:
$\exists A\in{\mathbb{R}}:\forall x\in{\mathcal{X}}:{\mathbf{E}}_{W(y|x)}{\left[{{\mathbf{N}}(y)}\right]}\leq A.$ (18)
We will call such channel an IDS channel. The IDS channel sequence is $D=D^{\infty}$, where $D_{n}$ is channel $W^{n}$, followed by a gluing operator on its output:
$D_{n}={\mathbf{g}_{n}}{\left({W^{n}}\right)},$ (19)
where $\mathbf{g}_{n}:{\overline{\mathcal{Y}}}^{n}\to{\overline{\mathcal{Y}}}$ glues together $n$ vectors over ${\mathcal{Y}}$. Such channel sequence $D$ is information-stable [3], so ${\mathbf{C}}(D)={\mathbf{I}}(D)$.
Proposition 6.
Any IDS channel (with finite expectation of output vector length) has finite output entropy for each input:
$\exists{\overline{H}}\in{\mathbb{R}}:\forall x\in{\mathcal{X}}:{\mathbf{H}}{\left[{W(y|x)}\right]}\leq{\overline{H}}.$ (20)
Proof.
Fix some $x_{0}\in{\mathcal{X}}$. Denote by $w_{y}=W(y|x_{0})$. The entropy is a series
${\mathbf{H}}{\left[{w_{y}}\right]}=\sum_{y\in{\overline{\mathcal{Y}}}:w_{y}>0}-w_{y}\log_{2}w_{y}.$ (21)
Since all the terms of the summation are positive, we only must prove that the sum is not +$\infty$+ ∞ to prove convergence. Denote by $p_{n}=\sum_{y\in{\mathcal{Y}}^{n}}w_{y}$ the probability that the length of output is $n$. From (18),
$\sum_{y\in{\overline{\mathcal{Y}}}}w_{y}{\mathbf{N}}(y)=\sum_{n=0}^{\infty}n\cdot\sum_{y\in{\mathcal{Y}}^{n}}w_{y}=\sum_{n=0}^{\infty}np_{n}\leq A.$
We obtained a constraint on $p_{n}$. Expression $\sum_{y\in{\mathcal{Y}}^{n}}-w_{y}\log_{2}w_{y}$ achieves its maximum (with fixed $p_{n}$) with uniform distribution within ${\mathcal{Y}}^{n}$, i.e., when $w_{y}=p_{n}/|{\mathcal{Y}}|^{n}$. Thus,
${\mathbf{H}}{\left[{w_{y}}\right]}\leq-\sum_{n=0}^{\infty}\sum_{y\in{\mathcal{Y}}^{n}}\frac{p_{n}}{|{\mathcal{Y}}|^{n}}\log_{2}\frac{p_{n}}{|{\mathcal{Y}}|^{n}}\leq{\mathbf{H}}{\left[{p_{n}}\right]}-A\cdot|{\mathcal{Y}}|.$
The second term is a finite constant. The first term is the entropy of distribution $p_{n}$, corresponding to the length of the output of $W$, which has finite expectation $\leq A$. The maximum possible entropy of discrete distribution over ${\mathbb{N}}\cup\left\{{0}\right\}$ with given (finite) mean is equal to the entropy of the geometric distribution, which is finite. ∎
4.2 The Markov-IDS channel sequence
Consider a discrete-time Markov chain with $s$ states ${\mathcal{S}}=\left\{{1,\dots,s}\right\}$. Let $G$ be an $s\times s$ matrix which defines the transition probabilities: $G_{\sigma,\tau}\in[0,1]$ is equal to the conditional probability of the next state being $\tau$ if the current state is $\sigma$.
Assume that each state $\sigma\in{\mathcal{S}}$ corresponds to a state channel $W_{1}[\sigma]:{\mathcal{X}}\rightsquigarrow{\overline{\mathcal{Y}}}$, which is an IDS channel. If the state channels have different output alphabets ${\mathcal{Y}}_{\sigma}$ for each $\sigma$, we can easily generalize this case by letting ${\mathcal{Y}}=\cup_{\sigma}{\mathcal{Y}}_{\sigma}$.
The finite-state Markov chain (FSMC) channel [5] works as follows:
  1. 1.
    The initial state $\sigma_{1}$ is given.
  2. 2.
    The $i$-th symbol $x_{i}\in{\mathcal{X}}$ is transmitted through channel $W_{1}[\sigma_{i}]$. The output is ${\overline{y}}_{i}\in{\overline{\mathcal{Y}}}$.
  3. 3.
    After the $i$-th symbol is transmitted, the system transits from state $\sigma_{i}$ to state $\sigma_{i+1}$ with probability $G_{\sigma_{i},\sigma_{i+1}}$.
  4. 4.
    For $n$ input symbols, one obtains ${\overline{y}}_{1}^{n}\in{\overline{\mathcal{Y}}}^{n}$.
Let $W_{*,1}[\sigma_{1}]=(W_{1,1}[\sigma_{1}],W_{2,1}[\sigma_{1}],\dots)$ be an FSMC channel sequence with initial state $\sigma_{1}$. Then
$W_{n,1}[\sigma_{1}](y_{1}^{n}|x_{1}^{n})=W_{1}[\sigma_{1}](y_{1}|x_{1})\cdot\sum_{\sigma_{2}^{n}\in{\mathcal{S}}^{n-1}}\prod_{i=2}^{n}G_{\sigma_{i-1},\sigma_{i}}W_{1}[\sigma_{i}](y_{i}|x_{i}).$ (22)
Also, introduce channel $W_{n}[\sigma]$ which outputs $\mathbf{g}_{n}({\overline{y}}_{1}^{n})\in{\overline{\mathcal{Y}}}$, where ${\overline{y}}_{1}^{n}$ is the output of $W_{n,1}[\sigma]$:
$W_{n}[\sigma]={\mathbf{g}_{n}}{\left({W_{n,1}[\sigma]}\right)}.$ (23)
Assume that the initial state of Markov chain is chosen according to distribution $\rho\in{\mathbb{D}}_{s}$. Then, the corresponding FSMC channel sequence is denoted by $V_{*,1}[\rho]$, $V_{n,1}[\rho]:{\mathcal{X}}^{n}\rightsquigarrow{\overline{\mathcal{Y}}}^{n}$:
$V_{n,1}[\rho](y|x)=\sum_{\sigma\in{\mathcal{S}}}\rho_{\sigma}\cdot W_{n,1}[\sigma](y|x).$ (24)
The Markov-IDS channel sequence $V[\rho]$, $V_{n}[\rho]={\mathbf{g}_{n}}{\left({V_{n,1}[\rho]}\right)}:{\mathcal{X}}^{n}\to{\overline{\mathcal{Y}}}$ is given by
$V_{n}[\rho](y|x)=\sum_{\sigma\in{\mathcal{S}}}\rho_{\sigma}\cdot W_{n}[\sigma](y|x).$ (25)
5 Capacity theorem for Markov-IDS channel
Throughout the section, we use notation of the channel sequences $W_{n,1}[\rho]$, $W_{n}[\rho]$, $V_{n,1}[\rho]$ and $V_{n}[\rho]$, $n\in{\mathbb{N}}$, as defined in (22), (23), (24) and (25).
5.1 Markov chain state distribution convergence
It is well-known (Theorem 8.9 in [19]), that if a Markov chain is aperiodic and irreducible, then
$\exists B\geq 0,0\leq P<1:\forall n\in{\mathbb{N}},\forall i,j\in{\mathcal{S}}:|G^{n}_{i,j}-\pi_{j}|\leq BP^{n}=B^{\prime}e^{-Cn}\stackrel{{\scriptstyle\triangle}}{{=}}\delta_{n}$ (26)
where $G^{n}_{i,j}=(G^{n})_{i,j}$, $\pi\in{\mathbb{D}}_{s}$ is the stationary distribution of Markov chain $G$: $\pi\cdot G=\pi$, constants $B,B^{\prime}\geq 0$, $10\leq P<$ 1 and $C>0$ depend only on the Markov chain.
For any starting distribution $\rho\in{\mathbb{D}}_{s}$ we have $\sum_{i}\rho_{i}=1$, and so
[…]
Introduce sequence $\varepsilon_{n}$, such that
$\delta_{n}=\varepsilon_{n}\min_{i}\pi_{i}.$ (28)
Then, the small (with $n\to\infty$) additive term can be substituted with small multiplicative term as $\pi_{i}+\delta_{n}\leq\pi_{i}(1+\varepsilon_{n})$. Combining this with (26), (27)
$\forall\rho\in{\mathbb{D}}_{s},\forall j:~{}$ $\pi_{j}(1-\varepsilon_{n})\leq(\rho G^{n})_{j}\leq\pi_{j}(1+\varepsilon_{n}),$
$\varepsilon_{n}\leq De^{-Cn},C=\operatorname{const},D=\operatorname{const},$
which also implies
$\pi_{j}(1-\varepsilon_{n})\leq G^{n}_{i,j}\leq\pi_{j}(1+\varepsilon_{n}).$
5.2 Main Theorems
Theorem 1 (Existence of i-capacity).
If a Markov chain $G$ of $s$ states, corresponding to Markov-IDS channel sequence $V$, is aperiodic and irreducible, then for any $\rho\in{\mathbb{D}}_{s}$ the i-capacity ${\mathbf{I}}(V[\rho])$ exists and does not depend on $\rho$. In other words, for any $\rho\in{\mathbb{D}}_{s}$, ${\mathbf{I}}(V[\rho])$ exists and
${\mathbf{I}}(V[\rho])={\mathbf{I}}(V[\pi]),$ (29)
where $\pi$ is the stationary distribution of the Markov chain.
Proof.
We now will prove that the limit ${\mathbf{I}}(V[\rho])=\lim_{n\to\infty}{\mathbf{I}}(V_{n}[\rho])/n$ exists. For that, we will use the stronger formulation [20] of Fekete’s lemma, which says that if function $u:{\mathbb{N}}\to{\mathbb{R}}$ is quasi-subadditive in the sense that
$\forall m,n\in{\mathbb{N}},n\leq m\leq 2n:u(m+n)\leq u(m)+u(n)+f(m+n),$ (30)
where term $f(m+n)$ is such that series $\sum_{n=1}^{\infty}f(n)/n^{2}$ converges to a finite number, then the limit $\lim_{n\to\infty}u(n)/n$ exists.
Denote by $V_{n_{1}|n_{2}|...|n_{t}}[\rho]:{\mathcal{X}}^{n}\rightsquigarrow{\overline{\mathcal{Y}}}^{t}$, $n=\sum_{i}n_{i}$ a channel, which works as a channel $V_{n,1}[\rho]$ combined with the merging operation. The merging operation concatenates the outputs inside $i$-th output sub-block for each $i$, where the corresponding $i$-th input sub-block has length $n_{i}$. So, the receiver knows the output, corresponding to each $n_{i}$-block, but does not know the exact output for each input symbol. Also, we write ${\widetilde{n}}_{i}$ instead of $n_{i}$ to denote the fact that the output, corresponding to the $i$-th input block is erased, i.e., substituted by the zero-length vector. Consider the case of $\rho=\pi$, i.e., the stationary distribution of the Markov chain (later we will show that the starting distribution does not influence the i-capacity). Then, the following sequence of inequalities holds  [^4] Here transitions $1$ and $2$ correspond to providing the receiver with the additional information. In transition $3$, let us assume the same input distribution $p\in{\mathbb{D}}[{\mathcal{X}}^{m+n}]$ for both $V_{m|l|n-l}[\pi]$ and $V_{m|{\widetilde{l}}|n-l}[\pi]$ and prove that ${\mathbf{I}}(V_{m|l|n-l}[\pi],p)\leq{\mathbf{I}}(V_{m|{\widetilde{l}}|n-l}[\pi],p)+l{\overline{H}}$. If it holds for any $p$, then we establish that the transition $3$ is also valid. Denote by $X_{1},X_{2},X_{3}$ the random variables corresponding to three input blocks of $m,l$ and $(n-l)$ symbols (they are the same for both channels since we take the same input distribution), and denote the random variables corresponding to the outputs by $Y_{1},Y_{2},Y_{3}$. Then, the amount of information about $X_{1}^{3}$ that we have lost by erasing $Y_{2}$ is given by ${\mathbf{I}}(V_{m|l|n-l}[\pi],p)-{\mathbf{I}}(V_{m|{\widetilde{l}}|n-l}[\pi],p)$ $=I(Y_{1}^{3};X_{1}^{3})-I(Y_{1},Y_{3};X_{1}^{3})=I(Y_{2};X_{1}^{3}|Y_{1},Y_{3})\leq H(Y_{2}|Y_{1},Y_{3})\leq H(Y_{2}),$ where $H(Y_{2})$ is the unconditional entropy of output corresponding to the second block; vector $Y_{2}$ consists of $l$ outputs $Y_{2}^{(i)}$ of component channels, glued together, $i=1\dots l$. So, $H(Y_{2})\leq H(Y_{2}^{(1)},Y_{2}^{(2)},\dots,Y_{2}^{(l)})\leq\sum_{i}H(Y_{2}^{(i)})$. Each $Y_{2}^{(i)}$ is an unconditioned output of a state channel. It can be expressed as $H(Y_{2}^{(i)})={\mathbf{E}}_{\sigma_{i},x_{i}}{\left[{H(Y_{2}^{(i)}|x_{i},\sigma_{i})}\right]}\leq\max_{\sigma,x}{\mathbf{H}}{\left[{W[\sigma](y|x)}\right]}=\overline{H},$ where distribution of $\sigma_{i}$ is given by initial state distribution $\rho$ and Markov chain transition probabilities, and distribution of $x_{i}$ is given by marginalization of distribution $p$ over all coordinates except $(m+i)$-th. The above implies that $H(Y_{2})\leq l{\overline{H}}$ and the inequality $3$ holds. In transition $4$ after block of $(n-l)$ symbols we further transmit block of $l$ symbols. This cannot decrease the mutual information, since we can just ignore last block of input and output. In transition $5$ the output of $(n-l)$-block is merged with the output of the $l$-block. We can additionally tell the receiver the exact values of output for last $l$-block, and then the receiver can know the border between the two outputs. By the same consideration as in inequality $3$, the conditional entropy about the last $l$ glued outputs, which we have partially lost by gluing, is not greater than unconditional entropy of $l$ unglued outputs, which is not greater than $l\overline{H}$ bits. This holds for each input distribution, so this holds for maxima over input distributions. :
${\mathbf{I}}(V_{m+n}[\pi])$ $\stackrel{{\scriptstyle 1}}{{\leq}}{\mathbf{I}}(V_{m|n}[\pi])\stackrel{{\scriptstyle 2}}{{\leq}}{\mathbf{I}}(V_{m|l|n-l}[\pi])\stackrel{{\scriptstyle 3}}{{\leq}}{\mathbf{I}}(V_{m|{\widetilde{l}}|n-l}[\pi])+l{\overline{H}}\stackrel{{\scriptstyle 4}}{{\leq}}{\mathbf{I}}(V_{m|{\widetilde{l}}|n-l|l}[\pi])+l{\overline{H}}$
$\stackrel{{\scriptstyle 5}}{{\leq}}{\mathbf{I}}(V_{m|{\widetilde{l}}|n}[\pi])+2l{\overline{H}}$ (31)
where $\overline{H}=\max_{\sigma\in{\mathcal{S}},x\in{\mathcal{X}}}{\mathbf{H}}{\left[{W[\sigma](y|x)}\right]}$ is the highest output entropy of a component channel (see Proposition 6).
Now, consider channel $V_{m|{\widetilde{l}}|n}[\rho]:{\mathcal{X}}^{m+l+n}\rightsquigarrow{\overline{\mathcal{Y}}}^{2}$. We will need auxiliary channel $W_{n,1}[\sigma\to\tau]:{\mathcal{X}}^{n}\to{\mathcal{S}}\times{\overline{\mathcal{Y}}}^{n}$, where $\tau$ is considered as a part of the channel’s output. The channel can be represented as channel $W_{n,1}[\sigma]$ which additionally outputs the state after transmission of the last symbol:
$W_{n,1}[\sigma_{1}\shortrightarrow\tau](y_{1}^{n}|x_{1}^{n})$ $=W_{1}[\sigma_{1}](y_{1}|x_{1})\cdot\sum_{\sigma_{2}^{n}\in{\mathcal{S}}^{n-1}}G_{\sigma_{n},\tau}\cdot\prod_{i=2}^{n}G_{\sigma_{i-1},\sigma_{i}}W_{1}[\sigma](y_{i}|x_{i}).$ (32)
Also, define the glued version of this channel as
$W_{n}[\sigma\shortrightarrow\tau]({\overline{y}}|x)$ $=\sum_{y\in\mathbf{g}_{n}^{-1}({\overline{y}})}W_{n,1}[\sigma\shortrightarrow\tau](y|x).$ (33)
The channel which does not output the final state and the channel which outputs the final state are related via
$W_{n}[\sigma]({\overline{y}}|x)=\sum_{\tau}W_{n}[\sigma\shortrightarrow\tau]({\overline{y}}|x),~{}W_{n,1}[\sigma]({\overline{y}}|x)=\sum_{\tau}W_{n,1}[\sigma\shortrightarrow\tau]({\overline{y}}|x).$
Now we can define the transition probabilities:
$V_{m|{\widetilde{l}}|n}[\pi](y_{1}^{2}|x_{1}^{3})\stackrel{{\scriptstyle\triangle}}{{=}}\sum_{\begin{subarray}{c}z_{1}\in\mathbf{g}^{-1}_{m}(y_{1})\\ z_{2}\in\mathbf{g}^{-1}_{n}(y_{2})\end{subarray}}\sum_{\sigma_{1}^{3}\in{\mathcal{S}}^{3}}\pi_{\sigma_{1}}W_{m}[\sigma_{1}\shortrightarrow\sigma_{2}]{\left({z_{1}|x_{1}}\right)}G^{l}_{\sigma_{2},\sigma_{3}}W_{n}[\sigma_{3}](z_{2}|x_{3})$ (36)
$\stackrel{{\scriptstyle 1}}{{\leq}}\sum_{z_{1}^{2},\sigma_{1}^{3}}~{}\pi_{\sigma_{1}}W_{m}[\sigma_{1}\shortrightarrow\sigma_{2}]{\left({z_{1}|x_{1}}\right)}\pi_{\sigma_{3}}(1+\varepsilon_{l})W_{n}[\sigma_{3}](z_{2}|x_{3})$
$\stackrel{{\scriptstyle 2}}{{\leq}}(1+\varepsilon_{l})\sum_{\sigma_{1},z_{1}}\pi_{\sigma_{1}}W_{m}[\sigma_{1}]{\left({z_{1}|x_{1}}\right)}\sum_{\sigma_{3},z_{2}}\pi_{\sigma_{3}}W_{n}[\sigma_{3}](z_{2}|x_{3})=(1+\varepsilon_{l})V_{m}[\pi](y_{1}|x_{1})V_{n}[\pi](y_{2}|x_{3})$ (37)
For brevity, introduce some shortcuts: let $x=x_{1}^{3}$, $y=y_{1}^{2}$. Since $V_{m|{\widetilde{l}}|n}[\pi](y_{1}^{2}|x_{1}^{3})$ does not depend on $x_{2}$, consider shortcut $U(y_{1}^{2}|x_{1}^{2})=V_{m|{\widetilde{l}}|n}[\pi](y_{1}^{2}|x_{1},x^{\prime},x_{2})$, where $x^{\prime}\in{\mathcal{X}}^{l}$ is some fixed vector. Also, consider shortcut $U^{\prime}(y|x)=V_{m}[\pi](y_{1}|x_{1})V_{n}[\pi](y_{2}|x_{2})$. Denote marginalized distribution ${\widetilde{p}}(x_{1},x_{2})=\sum_{x^{\prime}}p(x_{1},x^{\prime},x_{2})$. Rewriting (37) in shortcuts, one obtains $U(y|x)\leq(1+\varepsilon_{l})U^{\prime}(y|x)$, and one can upper-bound the mutual information by
${\mathbf{I}}(V_{m|{\widetilde{l}}|n}[\pi],p)={\mathbf{I}}(U,{\widetilde{p}})=\sum_{x,y}{\widetilde{p}}\circ U(x,y){\left[{\log_{2}U(y|x)-\log_{2}{\widetilde{p}}\circ U(*,y)}\right]}$
$\stackrel{{\scriptstyle 1}}{{\leq}}\sum_{x,y}(1-\varepsilon_{l}){\widetilde{p}}\!\circ\!U^{\prime}(x,y)\log_{2}{\left[{(1+\varepsilon_{l})U^{\prime}(y|x)}\right]}-(1+\varepsilon_{l}){\widetilde{p}}\circ U^{\prime}(x,y)\log_{2}{\left[{(1-\varepsilon_{l}){\widetilde{p}}\!\circ\!U^{\prime}(*,y)}\right]}$
$=(1-\varepsilon_{l}){\mathbf{E}}_{{\widetilde{p}}\circ U^{\prime}(x,y)}{\left[{\log_{2}U^{\prime}(y|x)+\log_{2}{\left({1+\varepsilon_{l}}\right)}}\right]}-(1+\varepsilon_{l}){\mathbf{E}}_{{\widetilde{p}}\circ U^{\prime}(x,y)}{\left[{\log_{2}{\widetilde{p}}\circ U^{\prime}(*,y)+\log_{2}{\left({1-\varepsilon_{l}}\right)}}\right]}$
$={\mathbf{I}}(U^{\prime},{\widetilde{p}})-\varepsilon_{l}{\mathbf{E}}_{{\widetilde{p}}\circ U^{\prime}(x,y)}{\left[{\log_{2}U^{\prime}(y|x)}\right]}-\varepsilon_{l}{\mathbf{E}}{\left[{\log_{2}{\widetilde{p}}\circ U^{\prime}(*,y)}\right]}+\Gamma_{l}$
$={\mathbf{I}}(U^{\prime},{\widetilde{p}})+\varepsilon_{l}{\mathbf{H}}_{{\widetilde{p}}\circ U^{\prime}(x,y)}{\left[{U^{\prime}(y|x)}\right]}+\varepsilon_{l}{\mathbf{H}}{\left[{{\widetilde{p}}\circ U^{\prime}(*,y)}\right]}+\Gamma_{l}$
$\leq{\mathbf{I}}(U^{\prime},{\widetilde{p}})+\varepsilon_{l}(m+n)\overline{H}+\varepsilon_{l}(m+n)\overline{H}+\Gamma_{l},$ (38)
where $\Gamma_{l}=(1-\varepsilon_{l})\log_{2}(1+\varepsilon_{l})-(1+\varepsilon_{l})\log_{2}(1-\varepsilon_{l})$. The transition in $\stackrel{{\scriptstyle 1}}{{\leq}}$ is done as follows. Each $\log$ in the first term adds a negative amount, so the multiplicative term before the first logarithm is lower-bounded. The logarithm itself is an increasing function, so its argument is upper-bounded. Note that the upper bound might turn the value of logarithm to a positive value. Then the whole term under $\sum$ sign is upper-bounded, because the original term was non-positive. The second term is taken with -- sign, so all bounds are the opposite.
Bringing together (31), (38), for $u_{n}={\mathbf{I}}(V_{n}[\pi])$ one obtains
$u_{m+n}\leq u_{m}+u_{n}+2\varepsilon_{l}(m+n)\overline{H}+2l\overline{H}+\Gamma_{l}$
Let $l=\lfloor{{\sqrt{m+n}}\rfloor}$ and denote by $f(m+n)=2\varepsilon_{l}\overline{H}+2l\overline{H}+\Gamma_{l}$. Then,
$0\leq f(n)\leq 2\varepsilon_{2\sqrt{2n}}+2\sqrt{2n}\cdot\overline{H}+\Gamma_{\lfloor{{\sqrt{2n}}\rfloor}}$
Obviously, $\varepsilon_{l}\to 0$ and $\Gamma_{l}\to 0$ with $l\to\infty$. Thus, $f(n)\in O(\sqrt{n})$, so the series $\sum_{n}f(n)/n^{2}$ converges. This implies the existence of $\lim_{n\to\infty}u_{n}/n={\mathbf{I}}(V[\pi])$.
We can show that ${\mathbf{I}}(V[\rho])={\mathbf{I}}(V[\pi])$ by the same trick with skipping $\lfloor{{\sqrt{n}}\rfloor}$ symbols at the beginning. ∎
To underline the fact that ${\mathbf{I}}(V[\rho])$ does not depend on $\rho$, we will always denote the i-capacity by ${\mathbf{I}}(V)$.
Theorem 2 (C-capacity equals i-capacity).
For any $\rho$, the coding capacity of $V[\rho]$ achieves the mutual information capacity of $V[\rho]$, and thus does not depend on $\rho$:
${\mathbf{C}}(V[\rho])={\mathbf{C}}(V)={\mathbf{I}}(V)$ (39)
Proof.
The scheme of the proof is presented in Fig. 1.
[…]
Figure 1: Scheme of the proof of Theorem 2
Step 1. Transition $V\to{\overline{V}}$.
Consider channel sequence ${\overline{V}}_{n}[\rho]=V_{l_{n}|m_{n}|l_{n}|...|m_{n}}[\rho]:{\mathcal{X}}^{n}\rightsquigarrow{\overline{\mathcal{Y}}}^{Q_{n}}$, for which the receiver will know the output corresponding to input subvectors of lengths $l_{n}$ and $m_{n}$, where
$m_{n}=\lfloor{{\sqrt{n}}\rfloor},~{}l_{n}=\lfloor{{\sqrt[4]{n}}\rfloor},~{}q_{n}=\bigg{\lfloor}\frac{n}{m_{n}+l_{n}}\bigg{\rfloor},~{}Q_{n}=2q_{n}$ (40)
and last $r_{n}=n-q_{n}(m_{n}+l_{n})$ input symbols are ignored. Note that with $n\to\infty$, the asymptotics of these variables are
$Q_{n}\sim q_{n}\sim m_{n}\in O(\sqrt{n}),~{}l_{n}\in O(\sqrt[4]{n}),~{}r_{n}\in O(n^{3/4}).$
Below, we will omit index ${}_{n}$ and write $l,m,q$ and $Q$ meaning that they depend on $n$ by (40).
Also, consider channel sequence ${\overline{V}}^{\prime}_{n}[\rho]$, which outputs the output of ${\overline{V}}_{n}[\rho]$ if the total length of output is $\geq 2An$, and otherwise outputs the output of $V_{n}[\rho]$. Formally, let
$\overline{\mathbf{g}}_{n}^{\circ}({\overline{y}}_{1}^{Q})=\begin{cases}{\overline{y}}_{1}^{Q}&\text{if }{\mathbf{N}}(\mathbf{g}_{Q}({\overline{y}}_{1}^{Q}))\leq 2An\\ \mathbf{g}_{Q}({\overline{y}}_{1}^{Q})&\text{if }{\mathbf{N}}(\mathbf{g}_{Q}({\overline{y}}_{1}^{Q}))>2An\end{cases}$ (41)
and then channel sequence ${\overline{V}}^{\prime}[\rho]$ is defined by ${\overline{V}}_{n}^{\prime}[\rho]=\overline{\mathbf{g}}_{n}^{\circ}({\overline{V}}_{n}[\rho])$. Note that, with $n\to\infty$, function $\overline{\mathbf{g}}_{n}^{\circ}(y)$ changes the input with probability 0 , i.e., $\Pr{\left[{\overline{\mathbf{g}}_{n}^{\circ}(y)\neq y}\right]}\to 0$. Moreover, the image of short sequences does not intersect the image of long sequences, so we can use Proposition 5 and obtain
${\mathbf{I}}({\overline{V}}^{\prime}[\rho])={\mathbf{I}}({\overline{V}}[\rho]),~{}{\mathbf{C}}({\overline{V}}^{\prime}[\rho])={\mathbf{C}}({\overline{V}}[\rho])$ (42)
Now let $\overline{\mathbf{g}}$ be a function which glues the vectors of vectors that remained unglued after $\overline{\mathbf{g}}^{\circ}$:
$\overline{\mathbf{g}}_{n}({\overline{Y}})=\begin{cases}\mathbf{g}_{Q}({\overline{Y}}),&\text{if }{\overline{Y}}\in{\overline{\mathcal{Y}}}^{Q}\wedge{\mathbf{N}}(\mathbf{g}_{Q}({\overline{Y}}))\leq 2An\\ {\overline{Y}},&\text{if }{\overline{Y}}\in{\overline{\mathcal{Y}}}\wedge{\mathbf{N}}({\overline{Y}})>2An\end{cases}$ (43)
On one hand, $V_{n}[\rho]=\overline{\mathbf{g}}_{n}({\overline{V}}^{\prime}[\rho])$. On the other hand, note that
$\Phi_{n}=\log_{2}\max_{{\overline{Y}}}\left|{\overline{\mathbf{g}}_{n}^{-1}({\overline{Y}})}\right|=\log_{2}\binom{2An+Q-1}{Q-1}\sim\log_{2}(2An)^{\sqrt{n}}\in O(\sqrt{n}\log n)\subset o(n).$
Thus, using by Proposition 3 we obtain ${\mathbf{I}}({\overline{V}}^{\prime}[\rho])={\mathbf{I}}(V)$ and ${\mathbf{C}}({\overline{V}}^{\prime}[\rho])={\mathbf{C}}(V[\rho])$, and together with (42) we obtain
${\mathbf{I}}({\overline{V}}[\rho])={\mathbf{I}}(V),~{}{\mathbf{C}}({\overline{V}}[\rho])={\mathbf{C}}(V[\rho]).$ (44)
Step 2. Transition ${\overline{V}}\to{\widetilde{V}}$.
Now, denote by ${\widetilde{V}}[\rho]$ the sequence of channels ${\widetilde{V}}_{n}[\rho]=V_{{\widetilde{l}}|m|{\widetilde{l}}|m|\dots}[\rho]$, where number of $l$- and $m$-blocks is $q$, and $l=l_{n}$, $m=m_{n}$, and $q=q_{n}$ are defined by (40). The difference between channels ${\widetilde{V}}_{n}[\rho]$ and ${\overline{V}}_{n}[\rho]$ is that each $l$-block is erased, so
${\mathbf{C}}({\widetilde{V}}[\rho])\leq{\mathbf{I}}({\widetilde{V}}[\rho])\leq{\mathbf{I}}({\overline{V}}[\rho]),{\mathbf{C}}({\widetilde{V}}[\rho])\leq{\mathbf{C}}({\overline{V}}[\rho])\leq{\mathbf{I}}({\overline{V}}[\rho])$ (45)
The total number of inputs, for which the outputs are erased, is $ql$. Assume some input distribution $p$ over ${\mathcal{X}}^{n}$. Then, the following random variables are defined: $X$ corresponds to the input; $Y$ is the erased output $l$-blocks; $Z$ is the output of ${\overline{V}}_{n}[\rho]$. Then,
${\mathbf{I}}({\overline{V}}_{n}[\rho],p)$ $=I(X;Y,Z)=I(X;Z)+I(X;Y|Z)={\mathbf{I}}({\widetilde{V}}_{n}[\rho],p)+I(X;Y|Z)$
$\leq{\mathbf{I}}({\widetilde{V}}_{n}[\rho],p)+H(Y)\leq{\mathbf{I}}({\widetilde{V}}_{n}[\rho],p)+lq{\overline{H}}$
Since $q\in O(\sqrt{n})$ and $l\in O(\sqrt[4]{n})$, the last term is in $o(n)$, which means that ${\mathbf{I}}({\overline{V}}[\rho])\leq{\mathbf{I}}({\widetilde{V}}[\rho])$. Recalling (45), one obtains ${\mathbf{I}}({\overline{V}}[\rho])={\mathbf{I}}({\widetilde{V}}[\rho])$.
Step 3. ${\widetilde{V}}$ is information-stable, and, thus, ${\mathbf{I}}({\widetilde{V}})={\mathbf{C}}({\widetilde{V}})$.
Similarly to (37), one can obtain
$(1-\varepsilon_{l})^{q}V^{q}_{m}[\pi](y_{\text{even}}|w)\leq{\widetilde{V}}_{n}[\rho](y_{1}^{Q}|x)\leq(1+\varepsilon_{l})^{q}V^{q}_{m}[\pi](y_{\text{even}}|w)$
where $w={\widetilde{f}}_{n}(x)$ is the input vector $x$ without $l$-blocks and the last incomplete block, i.e. ${\widetilde{f}}_{n}(x)=(x^{(1)},x^{(2)},\dots,x^{(q)})$ , $x^{(i)}=x_{i(m+l)-m+1}^{i(m+l)}$. Denote by ${\overline{p}}_{n}\in{\mathbb{D}}_{{\mathcal{X}}^{n^{\prime}}}$, $n^{\prime}=qm$ the marginalized distribution of ${\widetilde{f}}_{n}(x)$, given distribution $p_{n}$ over $x$:
${\overline{p}}_{n}(w)=\sum_{x:{\widetilde{f}}_{n}(x)=w}p_{n}(x)$
The fact that channel probabilities of ${\widetilde{V}}_{n}[\rho]$ and $V^{q}_{m}[\pi]$ are close implies the fact that mutual information density of ${\widetilde{V}}_{n}[\rho]$ and $V^{q}_{m}[\pi]$ are close too, since
$\mathbf{i}({\widetilde{V}}_{n}[\rho],p_{n},x,y)=\log_{2}\frac{{\widetilde{V}}_{n}[\rho](y|x)}{\sum_{x^{\prime}}p_{n}(x^{\prime}){\widetilde{V}}_{n}[\rho](y|x^{\prime})}\leq\log_{2}\frac{(1+\varepsilon_{l})^{q}\cdot V^{q}_{m}[\pi](y_{\text{even}}|w)}{(1-\varepsilon_{l})^{q}\cdot{\overline{p}}_{n}\!\circ\!V^{q}_{m}[\pi](*,y_{\text{even}})}$
$=\mathbf{i}(V^{q}_{m}[\pi],{\overline{p}}_{n},w,y_{\text{even}})+q\log_{2}\frac{1+\varepsilon_{l}}{1-\varepsilon_{l}}.$ (46)
The last term tends to zero with $n\to\infty$, since $\varepsilon_{l}\leq Ce^{-D\sqrt[4]{n}}$ decays exponentially fast and $q\in O(\sqrt{n})$. Similarly, $\mathbf{i}({\widetilde{V}}_{n}[\rho],p_{n},x,y)\geq\mathbf{i}(V^{q}_{m}[\pi],{\overline{p}}_{n},w,y_{\text{even}})-q\log_{2}\frac{1+\varepsilon_{l}}{1-\varepsilon_{l}}$, and thus
$\lim_{n\to\infty}{\mathbf{E}}{\left[{\mathbf{i}({\widetilde{V}}_{n}[\rho],p_{n},x,y)-\mathbf{i}(V^{q}_{m}[\pi],{\overline{p}}_{n},x,y_{\text{even}})}\right]}$
$=\lim_{n\to\infty}{\mathbf{E}}{\left[{\left|{\mathbf{i}({\widetilde{V}}_{n}[\rho],p_{n},x,y)-\mathbf{i}(V^{q}_{m}[\pi],{\overline{p}}_{n},x,y_{\text{even}})}\right|}\right]}=0.$
Now, let ${\overline{p}}_{n}$ be the direct product of $q$ distributions $p^{*}_{m}$ which are optimal for $V_{m}[\pi]$, i.e. let ${\overline{p}}_{n}(x_{1}^{q})=\prod_{i=1}^{q}p^{*}_{m}(x_{i})$ and let $p_{n}$ be any distribution such that ${\widetilde{f}}_{n}(p_{n})={\overline{p}}_{n}$. Then,
$\lim_{n\to\infty}{\mathbf{E}}{\left[{\frac{1}{n}\cdot\left|{\mathbf{i}({\widetilde{V}}_{n}[\rho],p_{n},x,y)-\sum_{i=1}^{q}\mathbf{i}(V_{m}[\pi],p^{*}_{m},x^{(i)},y_{\text{even}})}\right|}\right]}$
$=\lim_{n\to\infty}{\mathbf{E}}{\left[{\left|{\frac{\mathbf{i}({\widetilde{V}}_{n}[\rho],p_{n},x,y)}{n}-\frac{q}{n}\cdot{\mathbf{I}}(V_{m}[\pi])}\right|}\right]}=\lim_{n\to\infty}{\mathbf{E}}{\left[{\left|{\frac{\mathbf{i}({\widetilde{V}}_{n}[\rho],p_{n},x,y)}{n}-{\mathbf{I}}(V)}\right|}\right]}=0$ (47)
By (46), $\lim_{n\to\infty}{\mathbf{I}}({\widetilde{V}}_{n}[\rho],p_{n})/n={\mathbf{I}}(V)$, so $p_{n}$ allows channel sequence ${\widetilde{V}}_{n}[\rho]$ to have i-capacity equal to ${\mathbf{I}}(V)$ which means ${\mathbf{I}}({\widetilde{V}}_{n}[\rho])\geq{\mathbf{I}}(V)$. It can be shown [^5] Note that ${\widetilde{V}}_{n}[\rho]$ is obtained from $V^{q}_{m}[\pi]$ by deleting $o(n)$ component channel outputs, each of which has entropy upper-bounded by ${\overline{H}}$. Furthermore, inputs of $V_{m}^{q}[\pi]$ independently influence the output, so no distribution $p$ for ${\widetilde{V}}_{n}[\rho]$ can lead to ${\widetilde{p}}$ which has strictly larger m.i. capacity ${\mathbf{I}}(V^{q}_{m}[\pi],{\widetilde{p}})$, than independent maximization over distributions of each $x^{(i)}$. That said, we cannot lose more than $o(n)\cdot{\overline{H}}$ bits of information by transition from ${\widetilde{V}}_{n}[\rho]$ to $V^{q}_{m}[\pi]$. that ${\mathbf{I}}({\widetilde{V}}[\rho])={\mathbf{I}}(V)$. Thus, (47) implies
$\lim_{n\to\infty}{\mathbf{E}}{\left[{\left|{\frac{\mathbf{i}({\widetilde{V}}_{n}[\rho],p_{n},x,y)}{n}-{\mathbf{I}}({\widetilde{V}}_{n}[\rho])}\right|}\right]}=0,$
which is exactly the information stability condition for ${\widetilde{V}}[\rho]$, which by Proposition 1 implies that ${\mathbf{I}}({\widetilde{V}}[\rho])={\mathbf{C}}({\widetilde{V}}[\rho])$, and also, since ${\mathbf{I}}({\widetilde{V}}[\rho])={\mathbf{I}}(V)$, the m.i. capacity does not depend on $\rho$, so we can underline it by writing
${\mathbf{I}}({\widetilde{V}}[\rho])={\mathbf{C}}({\widetilde{V}}[\rho])={\mathbf{I}}({\widetilde{V}})={\mathbf{C}}({\widetilde{V}}).$ (48)
Recalling (45), one obtains ${\mathbf{C}}({\overline{V}})={\mathbf{I}}({\overline{V}})$, and by (44), ${\mathbf{C}}(V)={\mathbf{I}}(V)$.
6 Conclusion
In this paper, we have focused on channels with synchronization errors, for which the insertions/deletions are governed by a Markov chain. Specifically, we have shown that if the underlying Markov chain is stationary and ergodic, then the information capacity of the corresponding channel sequence exists and it is equal to the coding capacity. This result generalizes the classical result of Dobrushin on the existence of the Shannon capacity for channels with memoryless synchronization errors. To accomplish this goal, we generalized methods used in [2, 3] as separate independent propositions. These propositions state that one can apply a function sequence to a channel sequence, changing neither the coding capacity nor the mutual information capacity of the latter, if the function sequence has asymptotically small preimage or asymptotically small probability of changing the output. By applying such function sequences, one can bring the original channel sequence to another channel sequence, for which the capacity theorem or capacity bounds can be more easily derived, and one can be sure that the coding capacity / mutual information capacity is not changed under such transformation. The methodology may be useful for deriving capacity theorems and capacity bounds for other non-trivial channel sequences.